OpenStudy (anonymous):

how do you take the derivative of a summation? ie: question : find the derivative of the summation from 0 to infinity of (x^2n)/n!

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

You can recognize the function the series represents and then find the derivative of that function.

OpenStudy (anonymous):

Or, you can write out the first few terms of the given series, take the derivative term-by-term, and try to find a pattern in order to write the terms as a series.

OpenStudy (anonymous):

how do you recognize the function the series represents?

OpenStudy (anonymous):

i mean i guess it could represent e^2x?

OpenStudy (anonymous):

The power series for the exponential function \(f(x)=e^x\) is \[e^x=\sum_{n=0}^\infty\frac{x^n}{n!}\]

OpenStudy (anonymous):

Close, it's actually \(e^{x^2}\).

OpenStudy (anonymous):

oh so the derivative of e^x^2 is... ummm 2x e^x^2?

OpenStudy (anonymous):

Yes. And you can also write that as a power series.

OpenStudy (anonymous):

and thats the answer for the derivative of the summation?

OpenStudy (anonymous):

how do you do it as power series

OpenStudy (anonymous):

or take it term by term

OpenStudy (anonymous):

That depends on the scope of the question, but either series or function representation would probably be acceptable. As for writing the derivative as a power series: Since \(e^{x^2}=\displaystyle\sum_{n=0}^\infty\frac{x^{2n}}{n!},\) you have \[\begin{align*}2xe^{2x^2}&=2x\displaystyle\sum_{n=0}^\infty\frac{x^{2n}}{n!}\\ &=\displaystyle2\sum_{n=0}^\infty\frac{x^{2n+1}}{n!}\end{align*}\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

You're welcome.